in a school there are 560 children, there are 40 boys more than girls, how many girls are there?

Answer by macston(5194)

.
g=number of girls; b=number of boys=g+40
g+b=560 Substitute for b
g+g+40=560
2g=520
g=260 ANSWER: There are 260 girls
CHECK:
b=g+40=260+40=300
g+b=560
260+300=560
560=560